### v1

# A Modern Take on Visual Relationship Reasoning for Grasp Planning

Abstract
Interacting with real-world cluttered scenes pose several challenges to robotic agents that need to understand complex spatial dependencies among the observed objects to determine optimal pick sequences or efficient object retrieval strategies. Existing solutions typically manage simplified scenarios and focus on predicting pairwise object relationships following an initial object detection phase, but often overlook the global context or struggle with handling redundant and missing object relations. In this work, we present a modern take on visual relational reasoning for grasp planning. We introduce D3GD, a novel testbed that includes bin picking scenes with up to 35 objects from 97 distinct categories. Additionally, we propose D3G, a new end-to-end transformer-based dependency graph generation model that simultaneously detects objects and produces an adjacency matrix representing their spatial relationships. Recognizing the limitations of standard metrics, we employ the Average Precision of Relationships for the first time to evaluate model performance, conducting an extensive experimental benchmark. The obtained results establish our approach as the new state-of-the-art for this task, laying the foundation for future research in robotic manipulation. We publicly release the code and dataset at https://paolotron.github.io/d3g.github.io.
I Introduction
Every kitchen has a cluttered drawer where that specific tool you urgently need to complete a recipe is buried beneath various gadgets and stray items. Retrieving it requires carefully planning a sequence of grasps. Could robot assistants help with this task? As they progressively integrate into our daily lives, we would like to delegate increasingly complex procedures composed of multiple key steps to artificial agents. This requires comprehensive perception, understanding of highly unstructured environments and advanced planning skills. Kitchen drawers as well as industrial containers are just two of many possible examples of multi-object-stacking scenes requiring appropriate grasping sequences to prevent potential damages.
Previous research has framed this problem as visual manipulation relationship reasoning [1, 2, 3], highlighting as a primary challenge the inference of relationship among objects, which can be categorized as parent, child, or none. In this context, objects are nodes in a dependency graph, where parent nodes should be grasped after child nodes. Knowing the complete dependency graph enables the calculation of the shortest path to any object and facilitates planning the most efficient picking sequence to access it.
Existing algorithms address this task by either focusing solely on modeling relationships between pairs of objects, thereby neglecting the broader global context, or by inferring the graph from the entire image scene, which often leads to difficulties in handling redundant or missing edges [4, 1, 2]. Additionally, their approach treats object detection and manipulation relationship recognition as two distinct stages, with evaluation primarily concentrated on the latter. The experimental analyses usually focus on per-edge prediction, and typically run on scenes with a low (5 objects, [1]) or medium (19 objects, [5]) number of objects.
Figure 1: Manipulation relationship graphs provide essential information to plan grasping in complex scenes as those encountered in industrial bin picking. Our work introduces a new testbed for this task ➀. Moreover, we design a novel approach ➁ that considers both object detection and edge relationship prediction in an end-to-end process. Finally, we adopt a tailored evaluation metric ➂ to assess the method performance across different levels of difficulties and challenges.
With our work we propose a modern take on visual relationship reasoning for grasp planning and advance the field with the following contributions (see Fig. 1):
  • we introduce a new testbed based on [6]. We organize bin picking scenes in subsets of growing complexity based on the graph that outlines the objects’ dependencies, focusing on the number of nodes, connected subgraphs, and shortest paths from leaf to root. We name it D3GD: Dependency Graph Generation for Grasp planning Dataset, and contains scenes with up to 35 objects belonging to 97 different categories. It includes synthetic images and an additional real-world test split to study sim-to-real transfer;
  • we design a new end-to-end dependency graph generation model. We refer to it as D3G. Its objective is to jointly detect the objects and predict the adjacency matrix describing their spatial relationship. We formalize the task as a set prediction problem over the whole image and learn the edge features representing the graph relationships rather than relying on object localization and their feature post-hoc concatenation as done in previous work;
  • we assess the entire pipeline with a tailored metric. Standard evaluation strategies depend on implicit object scoring thresholds which may be difficult to set. Here we use for the first time the Average Precision of Relationships [7] to evaluate the generated dependency graph.
With a thorough experimental analysis we show how our approach outperforms several state-of-the art competitors in challenging experimental settings. Overall our work presents a novel benchmark suite that will be open-sourced to the community paving the way for future application-oriented research in robotic manipulation.
II Related Works
In this section we provide a brief overview of previous publications in research areas related to our work, starting from the basic components of object detection and graph learning. We also review existing literature on object relationship recognition and reasoning, considering applications to scene understanding and robotic manipulation.
Object Detection is an extensively studied task whose goal is to localize objects in a scene and recognize their category label. Detection approaches evolved from feature matching and part modeling [8], through two stages [9, 10] and one stage deep methods [11, 12], getting more recently to transformer-based techniques. The latter allows for global computation over a whole image and set prediction, leveraging attention between multiple objects as in the case of DETR [13]. The use of deformable attention has been shown to further boost such approaches [14].
Graph Learning is a powerful tool when studying complex interactions within structured data. Message passing based networks have been designed to leverage the sparsity and connectivity of graph structures, efficiently propagating information and capturing dependencies between nodes [15]. With the advent of self-attention based architectures [16], several efforts have been made to merge the graph neural network paradigm and the transformer one, obtaining competitive performance on graph benchmarks [17, 18, 19].
Relationship reasoning is an umbrella term that encompasses a variety of tasks where it is important to capture structural information from data and to model relations that can be simply geometric (spatial positioning) or semantic (e.g. how the action of a subject affects an object). Large datasets such as Visual Genome [20] and Open Images [7] paved the way for perception studies based on deep learning mainly dedicated to image captioning, with first solutions building on two-stage object detectors [21], and the latest ones based on transformers [22] with promising applications for the robotic field [23, 24].
In the area of decision-making processes and motion planning, generating the dependency graph is a pre-condition to grasping in cluttered scenes. Recent works have also included it as an intermediate [25, 26] or auxiliary objective when predicting robotic parallel grasp poses [27, 28, 29].
The first real-world visual manipulation relationship dataset (VMRD) was introduced in [1]. It is a collection of 5185 hand-annotated images depicting various configurations of objects in cluttered scenes each containing up to 5 object instances sampled from 31 categories. While valid, this dataset is limited in dimension and complexity. The authors also proposed a deep model including distinct stages for object detection and relationship prediction. The same high-level solution was then adopted by following approaches that used fully connected conditional random fields [4] and graph neural networks [2, 3]. These works included analyses on a larger relational grasp dataset named REGRAD [5] automatically collected in a virtual environment. It is composed of images annotated with manipulation relationship between objects in dense clutter and including multi-view perspectives [30]. Nonetheless, existing testbeds tend to focus on relatively simple scenarios and often lack comprehensive information on the experimental settings, making previous results difficult to reproduce [^1] Reproducible results on a dataset require knowledge of the pre-processing pipeline and access to baseline code and checkpoints. Unfortunately, most prior work in visual manipulation relationship reasoning lacks these resources. Specifically, in REGRAD, detailed analysis uncovered issues such as missing boxes, one-pixel masks, and inconsistent object classes. The lack of public code further impedes result replication. We overcome these limitations with our benchmark, re-implementing the competitor methods and releasing the code for fair comparison.. Lately with the release of the NVIDIA Isaac simulator [31] a new generation of highly realistic and accurately annotated datasets have been published [32, 6, 33], which provide the opportunity to scale up relational reasoning related tasks for robotics applications.
III Dataset
We define our experimental testbed D3GD to get closer to real-world conditions where robots are challenged with grasping objects from large containers. We build on MetaGraspNetV2 [6] which is composed of nearly 300K synthetic images and 2.3K real images depicting bin picking scenes composed by up to 35 object instances sampled from 97 classes. The dataset comes with both 2D and 3D modalities as well as various types of grasp and labels. For our use-case we consider only RGB images, the object bounding boxes, the grasping dependency graph and, for the baseline models that use them, the instance segmentation annotations.
We organize the dataset into different difficulty levels on the basis of three key metrics: number of objects, largest subgraph, and longest shortest path from leaf to root. When the number of objects in the scene increases, detecting them and reasoning on their relations may become more challenging due to occlusions. Thus, it is important to consider the scene density besides the object cardinality. We do that by measuring how many objects (nodes) are part of a connected subgraph within the whole dependency graph. Finally, we measure the longest shortest paths from a graph leaf (the first object to be picked) to the root (the last one). This means that we count the minimum number of objects that we need to remove before reaching an object at the bottom of the bin. As the dependency graph of a scene may have multiple leaves and roots, we annotate for each scene the longest of these paths which is a good indicator of how many layers of clutter the bin contains. Note that we are interpreting the dependency graph as a Directed Acyclic Graph (DAG) and we removed all the not-fitting scene samples (only $\sim$60) that, upon visual inspection, were clearly identified as errors produced during the synthetic dataset generation phase.
Overall, our D3GD is organized with difficulty sets as presented in Table I. For our study we disregard the trivial set and focus on the easy, medium, and hard ones (see examples in Figure 2). We split the dataset into train, eval, and test scenes with an 80:10:10 proportion.
TABLE I: Difficulty levels decomposition of our D3GD dataset. Each row describes the concurrent conditions that define a level set. For instance, a sample in the medium set is a scene with 6 or more objects, of which more than 40% are involved in connected subgraphs and showing at least 4 stacked objects.
Set # Objects Largest Longest # Samples
Subgraph Shortest Path
Trivial $<3$ $\geq 0$ $\geq 1$ 90798
Easy $\geq$ 3 $\geq 0$ $\geq 1$ 186998
Medium $\geq$ 6 $>$ 0.4 $\geq$ 4 13616
Hard $\geq$ 10 $>$ 0.5 $\geq$ 4 4884
Furthermore, we include in D3GD the real images of MetaGraspNetV2 as a cross-domain test set. Those samples (743 after filtering for missing and wrong annotations) are used to evaluate the generalization performance in a real-world scenario of models trained on synthetic data.
IV Method
In this section we present our new visual relationship reasoning model D3G. It is formalized as a transformer-based graph generation approach that predicts the graph nodes by detecting the objects in the scene and the edges representing the object spatial relationships (presence/absence and direction). We depart from commonly used GNNs architecture and propose a fusion between graph transformers and detection transformers. In this way all the components of the dependency graph can be obtained jointly. In contrast to previous works using two-stage pipelines [1, 2, 3], our end-to-end strategy avoids the use of feature-cropping layers that in high clutter environments tend to generate ambiguous features. Figure 3 presents a schematic illustration of the proposed approach and helps to visualize the model flow explained in the following.
Figure 2: Data samples from the three main difficulty levels of D3GD.
IV-A Detecting Objects as Graph Nodes
Given an input image of dimension $3\times H_{0}\times W_{0}$, we perform object detection as a direct set prediction problem by following DETR [13]. A CNN backbone extracts a compact feature representation of dimensionality $C\times H\times W$ (where $C=2048$, $H={H_{0}}/{32}$, $W={W_{0}}/{32}$) which is rescaled via a $1\times 1$ convolution to reduce the channel dimensionality. The resulting $d\times H\times W$ tensor is interpreted as a list of vectors ($d\times HW$). These are fed to a transformer encoder [16] together with their spatial positional encodings and the obtained output is passed to the transformer decoder which also receives $N$ object queries (tensor of dimensionality $N\times F_{0}$, initially set to zero). The latter are transformed by the decoder in new embeddings that maintain the original dimension and are then processed by a feed-forward network into box coordinates and class labels. By utilizing self-attention and encoder-decoder attention mechanisms, the model globally reasons about all the objects at once, considering their pairwise relations while leveraging the entire image as context.
IV-B Predicting Object Relations as Graph Edges
To elaborate the graph edges, we provide the $N$ embeddings produced by the decoder as input to a graph transformer inspired by [18]. The node embeddings go through three different mappings operated by the matrices $M_{h}$, $M_{\epsilon 1}$, $M_{\epsilon 2}$ that work as adapters (each of dim. $F_{0}\times F$). We indicate their outputs respectively as $h_{0}$, $\epsilon_{1}$, $\epsilon_{2}$ (each of dim. $N\times F$). The last two are pairwise concatenated and then rescaled by halving the dimensionality via a further mapping $M_{e}$, to obtain $e_{0}=M_{e}(\epsilon_{1}\oplus\epsilon_{2})$. Both $h_{0}$ (dim. $N\times F$) and $e_{0}$ (dim. $N\times N\times F$) enter the first graph transformer layer and are progressively refined.
Figure 3: Schematic illustration of our D3G. We can identify three main rows in this figure. The top row describes the transformer architecture composed of encoder and decoder. The former takes as input scene features and provides as output updated representations which are fed to the decoder together with a fixed number of object queries. The query features obtained from the decoder are filtered and matched to the ground truth objects in the scene to detect them during training. The middle row shows how the same object query representations define the edge features via pairwise concatenation and mappings, before being refined by several graph transformer layers. Finally, a binary classifier predicts from the edge features whether two objects in the scene are connected by a spatial dependency relation. The bottom row zooms inside a layer of the graph transformer assuming only one head for visual clarity.
Each layer of our graph transformer operates similarly to the classical multi-head attention [16] but integrates edge features in its pipeline. A layer is characterized by an embedding of dimension $F$ and by $S$ heads. We conveniently define the dimension of the output of each head as $z=F/S$. We denote by $s=1,\ldots,S$ the index that runs over the $S$ heads, and by $l=1,\ldots,L$ the index over the $L$ layers.
A layer takes as input the node features $h_{l-1}$ and edge features $e_{l-1}$ and elaborates them through three learned projection matrices (each of dim. $F\times z$), query $Q_{l,s}$, key $K_{l,s}$ and edge $E_{l,s}$, to produce a pseudo-attention matrix $\hat{A}_{l,s}$ (dim. $N\times N\times z$) whose $i,j$ elements are obtained as
$\hat{A}_{l,s}^{ij}=\left(\frac{(Q_{l,s}\cdot h_{l-1})\cdot{(K_{l,s}\cdot h_{l-1})}^{\intercal}}{\sqrt{z}}\right)\cdot(E_{l,s}\cdot e_{l-1}^{ij})~{}.$ (1)
By summing over the third feature dimension ($z$) and applying a column-wise softmax operation we get the attention matrix in standard format (dim. $N\times N$):
$A_{l,s}=\text{softmax}(\text{sum}_{z}(\hat{A}_{l,s}))~{}.$ (2)
Finally, the layer output is composed of the updated node and edge features obtained respectively as:
$h_{l}$ $=O_{h}^{l}\|_{s=1}^{S}(A_{l,s}\cdot V_{l,s}\cdot h_{l-1})$ (3)
$e_{l}$ $=O_{e}^{l}\|_{s=1}^{S}\hat{A}_{l,s}~{}.$ (4)
[…]
IV-C Loss Function
To train our D3G we need to evaluate jointly its object detection and object relationship predictions, and progressively optimize them. The loss function, which measures both aspects, must accommodate the model’s output of a fixed set of $N$ bounding boxes, which exceeds the typical number of objects $P<N$ in an image. Thus, we need to search for the optimal match between the predicted and the ground truth objects.
We define as $y$ and $\hat{y}$ the set of ground truth and predicted objects. They are both of dimension $N$, with the former containing $P-N$ positions padded with $\emptyset$ (no object). More precisely, it holds $y_{i}=(c_{i},b_{i})$, where the class label $c_{i}$ may be $\emptyset$, and the ground truth box $b_{i}\in[0,1]^{4}$ is composed of the center coordinates as well as its height and width relative to the image size. By following [13] and using the Hungarian algorithm, we search for the optimal bipartite matching $\hat{\sigma}$ from all possible permutations of the $N$ elements $\sigma\in\mathfrak{G}_{N}$ with
$\hat{\sigma}={\arg\min}_{\sigma\in\mathfrak{G}_{N}}\sum_{i}^{N}\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})~{},$ (5)
$\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})=$ $-\mathds{1}_{\{c_{i}\neq\emptyset\}}\hat{p}_{\sigma(i)}(c_{i})$
$+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\text{box}}(b_{i},\hat{b}_{\sigma(i)})~{}.$ (6)
Here $\hat{p}_{\sigma(i)}(c_{i})$ is the probability of class $c_{i}$ for the prediction with index $\sigma(i)$, and $\hat{b}_{\sigma(i)}$ is the related bounding box. The loss component $\mathcal{L}_{\text{box}}$ is a linear combination of a $\ell_{1}$ loss and a generalized scale-invariant IoU loss [34].
Once the matching is obtained, we leverage the optimal assignment $\hat{\sigma}$ to calculate the node loss that measures the accuracy in detecting the objects that define the nodes of the dependency graph, and the relationship loss that evaluates the accuracy in the edge prediction. The first is a linear combination of a negative log-likelihood for class prediction and box loss [13]:
$\mathcal{L}_{node}(y,\hat{y})=\sum_{i=1}^{N}\left[-\log\hat{p}_{\hat{\sigma}(i)}(c_{i})+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\text{box}}(b_{i},\hat{b}_{\hat{\sigma}(i)})\right],$ (7)
it differs from $\mathcal{L}_{match}$ by the use of the log-probability in the first term and by calculating a cost for all the classes, including the assignment to a void object with class $\emptyset$.
Regarding the edge relationship, we start by introducing the rectangular selection matrix $T$ of dimensions $N\times P$ whose elements are defined as follows:
$T_{ij}=\mathds{1}_{\{i=\hat{\sigma}(j)\}}$ (8)
where $i$ and $j$ run respectively over rows and columns (see Figure 4). We use it to focus only on the identified ground truth objects that represent the nodes of the dependency graph and evaluate the related edges via the predicted adjacency matrix:
$\hat{G}=T^{\intercal}\cdot e_{pred}\cdot T~{}.$ (9)
Finally, we calculate the relationship loss using a simple binary cross entropy that indicates whether we correctly identify the presence or absence of an edge:
$\mathcal{L}_{rel}=\sum_{i}^{P}\sum_{j}^{P}\mathcal{L}_{bce}(G_{ij},\hat{G}_{ij})~{},$ (10)
where the binary ground truth matrix $G$ represents the object relationships and their direction ($G_{ij}\neq G_{ji}$). Overall, the total loss formulation for our D3G is $\mathcal{L}_{tot}=\mathcal{L}_{node}+\mathcal{L}_{rel}$.
Figure 4: Visualization of the loss calculation procedure. The optimal bipartite matching $\hat{\sigma}$ identifies the correspondence between the predictions ($N=5$) and the ground truth objects ($P=3$). Specifically, in this example $y=[1,2,3,0,0]$, where the three objects in order are Spray 1, Spray 2 and Pringles. The prediction is $\hat{y}=[3,0,2,0,1]$ and $\hat{\sigma}=[5,3,1,2,4]$. The matrix $T$ is obtained from $\hat{\sigma}$ starting from an $N\times P$ zero matrix and filling with a 1 only those position for which Eq. 8 hold, e.g. $i=1,j=3$ $\hat{\sigma}(j=3)=1$. The predicted adjacency matrix $\hat{G}$ defined in Eq. 9 is finally compared with the ground truth one.
V Metrics
The task of visual relationship reasoning for grasp planning is usually evaluated by adopting metrics from the detection literature as the (mAP) that is the mean of the Average Precision over all the object categories at a given Intersection over Union ($IoU>0.5$). For each class, the AP is the average precision when the value of recall increases from 0 to 1 (area under the precision vs recall curve). Indeed, this metric does not depend on the absolute probability value assigned to each bounding box for a certain class, rather it sweeps over all the bounding boxes assigned to that class sorted by probability (only the relative score matters).
Since the exact probability value is highly sensitive to the calibration process, disregarding it allows for a more objective comparison among different approaches. The Object triplet Recall (OR) and Precision (OP) are computed on object pairs, considering the relational triple $(obj_{i},r_{ij},obj_{j})$ true when all elements are correctly predicted. Specifically, the edge between two objects may have three possible labels $r:(in,out,no\_rel)$. OR and PR are the average recall and precision of all the three kinds of manipulation relationships.
These metrics depend on an object prediction score threshold, used to determine whether a bounding box contains an object at an IoU greater than 0.5. This threshold is typically tuned on the validation set which might not be available or be very dataset-specific, leading to generalization challenges. To address this limitation, we propose an alternative metric.
TABLE II: Results on the D3GD synthetic set. Best results per column highlighted in bold and second best underlined.
Model Detector Base Synth Easy Synth Medium Synth Hard
mAP $\text{AP}_{\text{rel}}$ OR OP mAP $\text{AP}_{\text{rel}}$ OR OP mAP $\text{AP}_{\text{rel}}$ OR OP
VRMN [1] Mask-RCNN [10] 86.96 73.29 88.76 86.12 75.20 54.97 70.22 71.86 67.36 42.49 62.26 71.58
GGNN [3] 86.31 70.51 87.41 84.15 74.19 51.43 68.04 69.39 66.02 37.52 59.11 68.58
GVRMN [2] 86.39 57.22 83.27 80.58 74.03 34.89 62.63 63.30 66.05 21.54 54.34 60.50
EGTR [22] Def. DETR [14] 84.23 52.47 79.28 78.20 75.37 33.13 61.33 70.11 66.66 22.72 48.07 73.38
D3G (ours) DETR [13] 83.22 76.27 89.57 88.64 70.92 58.65 72.86 71.58 60.25 43.99 63.71 57.85
Def. DETR [14] 87.68 76.96 87.42 90.66 74.84 60.81 66.91 77.98 65.92 46.45 51.69 85.18
[…]
The $\text{AP}_{\text{rel}}$ was originally introduced for visual relationships detection on Open Images [7], however it provided very pessimistic results and was soon replaced by other metrics. The cause of this behavior should be traced to the extensive variety of scenes and multiple valid semantic relations that may exist between the same pairs of objects, making comprehensive labeling extremely challenging. This is not the case for our task and data collection, where the object relationships can be exhaustively annotated.
VI Experiments
VI-A Reference Baselines
We compare our D3G with three previous approaches all composed of separate object detection and relationship reasoning stages. VMRN [1] originally exploited a Single Shot Detector (SSD [12]). The obtained bounding box and related convolutional features are provided as input to an Object Pairing Pooling Layer to classify manipulation relationships. Both GVRMN  [2] and GGNN [3] build on Faster-RCNN [9] including an object pair pooling to provide paired features to a graph neural network. Specifically, GVRMN includes a relation filtering network that prunes unrelated object pairs from the full relational graph and generates a compact one. Instead, GGNN leverages a gate recurrent unit that improves the long-term information passing of nodes in the graph, facilitating the integration of different nodes in the graph structure. For a fair comparison, we follow the best practice proposed in [6], considering Mask-RCNN with ResNet50 backbone as detector for all these three approaches and adding the respective network heads. We highlight that Mask-RCNN exploits instance segmentation as an auxiliary objective and thus needs the related data annotations.
We also add a fourth reference baseline from the scene graph generation literature. The recent EGTR [22] presented a lightweight object relation extractor from the decoder of a deformable DETR detector [14]. Query and key representation from each self-attention layer are concatenated pairwise and, together with the last hidden representation of the object queries, are fed to a gated sum to aggregate information before predicting the relation with a shallow head. A multi-task learning strategy allows to predict jointly the object location and category as well as the relationship label. As we will discuss in the result subsection, the graph transformer layer exploited by our D3G allows to better manage complex dependency relations for grasp planning. We evaluate D3G when built on DETR and on deformable DETR from a ResNet50 backbone.
VI-B Training Procedure Details and Experimental Setup
We train all the considered methods by first focusing on the detection modules and reaching convergence. This process takes around 12 epochs for Mask-RCNN and 18 epochs for the DETR based architectures. The relational reasoning components are then trained while further fine-tuning detection with a lower learning rate for 6 epochs: the high number of repeated scenes in different viewpoints leads to relatively fast convergence. For all phases we use AdamW optimizer [35]. We tune all the hyperparameters by monitoring the mAP and the $\text{AP}_{rel}$ on the validation set. Finally, we report all results on the three synthetic test split at different difficulty levels and the real set when using the chosen hyperparameters and training on the combined set of both training and validation data.
We use random flip, brightness, contrast, and saturation data augmentations. Notably, we observed that color augmentations are particularly effective in narrowing the performance gap between synthetic and real datasets. We conduct all experiments on a single GPU NVIDIA A100(40GB).
VI-C Results on D3GD
[…]
TABLE III: Results on the D3GD Real set. Best results per column highlighted in bold and second best underlined.
Model Detector Base Synth-to-Real
mAP $\text{AP}_{\text{rel}}$ OR OP
VMRN [1] Mask-RCNN [10] 45.04 13.36 49.70 10.20
GVRMN [2] 43.76 7.62 49.30 10.30
GGNN [3] 44.50 10.55 39.03 7.11
EGTR [22] Def. DETR [14] 56.36 15.11 56.25 48.88
D3G (ours) DETR [13] 55.93 24.44 41.11 51.36
Def. DETR [14] 56.14 22.38 33.70 38.90
[…]
Figure 5: Qualitative example of VRMN vs D3G. The bounding boxes show the detection predictions, while the graphs on the right present the related dependency relations. The object localization is almost identical and correct in the two cases. Regarding the graph, VRMN shows a wrong relationship between objects 0 and 2: to represent it the model focuses on pair pooling features obtained from the union area of the two bounding boxes that include an intruding object causing confusion. Thanks to the concept of object queries and the feature refinement provided by the transformer network, D3G offers a more unambiguous way of encoding relationships and infers the right dependency graph.
VI-D Analysis of Design Choices
In designing our D3G, we made specific choices regarding how the representations of two nodes (objects) are initially combined to define an edge (relationship) and the composition of the layers in the graph transformer module. Here we discuss these choices in detail, showing the impact of alternative options.
TABLE IV: Performance differences when changing the initialization technique for the edge representation.
$e_{0}$ init. strategy S.Easy S.Medium S.Hard Real
Hadamard product 76.20 58.93 44.23 24.09
Sum 75.96 59.51 43.61 21.45
Difference 76.42 59.81 44.41 23.59
Pairwise Concat. 76.27 58.65 43.99 24.44
TABLE V: $\text{AP}_{\text{rel}}$ performance when changing the graph head size $S$ and the number of layers $L$.
Graph Transf. # Params S.Easy S.Medium S.Hard Real
S=2 L=1 0.921M 76.26 58.08 44.30 25.32
L=2 1.84M 76.77 58.65 43.99 24.44
L=4 3.68M 76.06 59.32 44.25 25.37
S=4 L=1 0.921M 76.40 60.00 45.27 23.18
L=2 1.84M 76.10 54.84 44.17 24.66
L=4 3.68M 73.38 57.68 50.31 24.32
To define $e_{0}$, D3G exploits the pairwise concatenation of $\epsilon_{1}$ and $\epsilon_{2}$. We tested three other strategies: Hadamard product, Sum, and Difference. The results, presented in Table IV, show minimal variability between the various methods the best are pairwise concatenation and pairwise differences, we choose the former as a good tradeoff between synthetic and real performance.
The graph transformer, which is at the heart of our D3G, can come in different sizes depending on the number of heads $S$ in the multi-head self-attention of each layer, and on the number of layers $L$. We adopted $S=2$, $L=2$, but we can evaluate a smaller ($L=1$) and larger ($L=4$) version obtained by changing the number of layers. We also test the effect of increasing the number of heads to $S=4$. The obtained results in Table V indicate that on the easy set the results remain almost stable except when the number of heads and layers are too high ($S=4,L=4$) which leads to overfitting. On the medium and hard sets, increasing the number of heads can be beneficial but may require a significant increase in model parameters (e.g. $S=4,L=4$ for the hard case). Finally, the results on the real data indicate that it is preferable to maintain a small number of heads ($S=2$), and reducing the number of layers may even provide some further advantage ($L=1$ vs $L=2$).
VII Conclusions
With this work we shed new light on the task of visual manipulation relationship reasoning. We introduce D3GD, a new testbed that allows to scale up to more complex scenes than those usually considered in previous studies. Moreover, we move from the standard cascade strategy that requires first to detect the objects, usually with a two-stage detector, and then adopting pairwise pooling to combine the features extracted from the object bounding boxes. We leverage a transformer-based detector and graph layers to obtain our D3G end-to-end model that jointly predicts object location, class label, and pair relationships. Additionally, we point out that threshold-free evaluation metrics should be preferred for their reliability, especially in all those real-world cases in which a validation set to tune an ad-hoc threshold is not available. Thus, we propose to gauge and compare the prediction abilities of different approaches through the Average Precision of Relationships. We publicly release all the setting details of our testbed and experimental analysis, together with the code suite to foster reproducibility.
While the performance of our model on the proposed benchmark is remarkable there is still space for improvement. In particular, the graph transformer module handles a fully connected graph to model interactions among object queries, which may require considerable memory resources during training. This limitation can be addressed in future work by implementing subquadratic attention mechanisms by taking inspiration from [36]. Additionally, we plan to extend the task to more advanced reasoning levels, incorporating multimodal signals such as 3D cues and language. Overall we expect that the presented advancements will spur further exploration into dependency graph generation for robotic manipulation.
ACKNOWLEDGMENT
This work was carried out within the FAIR - Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013.